# LlamaTouch: A Faithful and Scalable Testbed for Mobile UI Automation Task Evaluation

### v1

@@ 4. Evaluation
4.1. Dataset
Table 3. LlamaTouch’s dataset statistics.
Task Category # Tasks # Apps Avg. Task Steps
Tasks from AITW (Rawles et al., 2023) 102 26 7.35 (2-19)
LlamaTouch Generated 393 46 5.67 (3-42)
Total 495 57 7.01 (2-42)
LlamaTouch uses a combination of tasks in the previous AITW dataset (Rawles et al., 2023) and self-constructed ones for evaluation. Table 3 shows the statistics of this dataset. We sampled 102 tasks from AITW (Rawles et al., 2023), one of the most popular datasets for mobile device control. By observing the limited number of evaluated applications, we generated 395 tasks in 46 applications. Overall, this dataset includes 495 tasks, covering 57 popular Android applications with different difficulty levels (measured by the number of steps to finish a task). Every single data entry consists of three types of data.
  • A task description in natural language, e.g., empty the shopping cart on bestbuy.
  • A sequence of raw UI representations when recording these tasks. Each raw UI representation contains a screenshot, VH, activity name, and action performed on the current screen.
  • The annotated essential states by human annotators.
@@ 4.2. Experiment Setup
Mobile environment. LlamaTouch primarily utilizes an x86-64 Android emulator (Google, 2023) as the mobile environment for task execution. The Android emulator is configured with Android 12 (API level 31) with Google Play Services deploying on a Linux server with Ubuntu 18.04 OS. Tasks involving applications that are not compatible with the Android emulator (e.g., Snapchat) will be tested on a real Google Pixel 5 device with the Android 14 OS.
@@
  • AutoUI (Zhan and Zhang, 2023) that employs a multimodal transformer model based on BLIP-2 (Li et al., 2023) and t5 (Raffel et al., 2020) with its variants as the brain for decision making. It takes pixel-level screenshots and task descriptions as the input.
  • AutoDroid (Wen et al., 2023) that uses LLMs for device control. It first takes textual screen VH and simplifies it to explicit, readable HTML representation. Then, the simplified HTML representation will be processed by LLMs and generate the corresponding action. We selected GPT-4 (gpt-4-0125-preview) (Achiam et al., 2023) as the backend of AutoDroid.
  • AppAgent (Yang et al., 2023) that utilizes GPT-4V (OpenAI, 2023) to comprehend the screenshots of mobile devices and then dispatch controlling commands to LlamaTouch. In our evaluation, we used the version AppAgent without document pre-exploration for simplicity.
  • CoCo-Agent (Ma et al., 2024) that uses LLaVA (Liu et al., 2023) as the brain. It is trained on a subset of the original AITW dataset.
All actions generated by these agents are applied to LlamaTouch’s Android environment through one adaption layer.
@@
  • Step-wise action match is widely-utilized in evaluating mobile agents on well-established datasets (Rawles et al., 2023; Sun et al., 2022; Wen et al., 2023; Zhan and Zhang, 2023). It compares the agent-generated action sequences in real-world environments with ground-truth action sequences in the dataset. When two action sequences are identical, the task is treated as completed. We utilize the action match algorithm in AITW (Rawles et al., 2023): two actions are matched only if they have identical action types and parameters.
@@
Metrics. We primarily compare the (1) end-to-end task completion rate (TCR) of different agents and (2) the accuracy of different evaluation approaches, taking human validation as the ground-truth result. In step-wise action match and LCS-based action match, a task is completed only when the two action sequences are identical or the ground-truth one serves as the subsequence of the agent-generated action sequence, respectively. In LlamaTouch, a task is treated as completed when the agent execution trace passes through all essential states. For accuracy, a task is accurately evaluated only if the evaluation result matches the human validation result. The accuracy of an evaluation approach denotes the proportion of tasks correctly evaluated out of all tasks.
@@ 4.3. Task Completion Rate and Accuracy
To demonstrate LlamaTouch’s effectiveness in addressing the issue of false negative results that pervasively occurred in previous evaluation methods, we focused on tasks considered successfully completed by human evaluation, excluding all incomplete tasks. Table 5 displays the number of tasks completed by agents and the evaluation accuracy for these tasks. Overall, agents successfully completed 30 tasks on average. Among these tasks, both step-wise action match and LCS-based action match achieve no more than 0.1% evaluation accuracy. This indicates they are unable to faithfully evaluate tasks executed in real-world environments using static datasets. Notably, LlamaTouch exhibits an average accuracy of 79% in validating these task execution traces, significantly reducing the percentage of false negative cases observed with other evaluation methods. The findings uncover LlamaTouch’s proficiency in evaluating UI automation tasks under real-world settings.
@@
$4.$ Ablation Study
@@
Evaluation design All tasks AITW tasks LlamaTouch tasks
@@
LlamaTouch achieves high accuracy when evaluating agent execution traces in real-world environments, attributing to the annotation and implementation of different essential state primitives. In this section, we evaluate the effectiveness of two types of match design in LlamaTouch: fuzzy match and exact match. We aggregate the execution traces of all mobile agents evaluated in $\S$§4.3. As for evaluation implementation, we disable exact match and fuzzy match separately, and then enable different primitives once per time. We present the results on (1) all tasks; (2) tasks in AITW (Rawles et al., 2023); and (3) tasks generated in LlamaTouch. The results of task completion rate and accuracy are shown in Table 6.
Exact match ensures key information in two screens are identical. As shown in the results, exact match contributes a significant portion of evaluation accuracy improvement. For example, without all exact match primitives, the evaluation accuracy of LlamaTouch on all tasks drops from 95% to 19%. Among all exact match primitives, activity match highly improves the validation accuracy. This is not only attributed to its simple yet efficient ability to locate functional application screens, but also the large portion of annotated activity primitives, i.e., 35% out of all primitives in our dataset. UI component match is also necessary for faithful evaluation in LlamaTouch, as it is typically with screen location primitives such as activity to detect critical information within the matched screen. Action and system state exact match only improve a little accuracy as only a few tasks are annotated with these primitives.
Fuzzy match does not show notable improvement in evaluation accuracy, although there are 30 screen-level fuzzy match primitives and 52 textbox fuzzy match primitives in the dataset. Seldom of tasks with these primitives can be completed by the four agents we evaluated, leading to a slight improvement. Although primitives about fuzzy match are not well explored, they play an important role in dealing with screen content or UI layout dynamics in the real world. With more performant mobile agents in the future, they will be further explored and evaluated.
@@ 4.5. Absolute Capabilities of Mobile Agents
In this section, we present the absolute capabilities of different mobile agents in mobile UI automation tasks. We first categorize tasks according to their sources: AITW and our self-constructed dataset. This aims to reveal whether an agent that has previously learned on a well-established dataset has the capability to adopt new applications/tasks. Second, we categorize tasks according to their difficulty, measured by the number of steps required to complete a task in the ground-truth dataset. Tasks are further divided into three difficulty levels: easy (steps $\leq$ 4), medium (4 <<< steps $\leq$ 8), and high (step >>> 8).
Table 7. The comparison of task completion rate categorized by task sources (e.g., from the previous dataset AITW or from LlamaTouch’s generated dataset).
@@
Performance among datasets. Among the four evaluated agents, AutoUI and CoCo-Agent are previously trained on AITW. AutoDroid and AppAgent directly invoke GPT-4 and GPT-4V, respectively. We separately show end-to-end TCR of all agents on tasks in AITW and LlamaTouch’s generated tasks to show their performance generalization to new scenarios (i.e., new tasks and new applications). Results in Table 7 show that AutoDroid and AppAgent also show higher TCR in AITW tasks than LlamaTouch’s generated tasks. This mainly attributes to the task complexity gap between the two datasets. AutoUI and CoCo-Agent achieve 12.75% TCR on tasks in AITW. However, for generated tasks in LlamaTouch, both of them operate poorly: with only 2.29% and 2.31% TCR, respectively. The TCR gap between datasets reveals their lack of capability to adapt to unseen tasks before. Considering massive real-world applications, mobile agents with strong generalization capabilities to unseen scenarios are more competitive.
Performance under different task complexity. We categorize tasks into three difficulty levels according to the step number required to complete the task in the datasets. The results are shown in Table 8. Generally, mobile agents can better complete simple tasks that require fewer steps. There is a significant drop in TCR when task complexity increases. AutoDroid outperforms the other three agents for tasks under all difficulties. AppAgent achieves a similar TCR among all tasks as AutoDroid. We deem that the knowledgable GPT-4/4V can better interpret task descriptions in natural language and pixel-level or textual UI representations. Therefore, agents based on GPT-4/4V achieve high TCR. However, there is still huge space for mobile agents to improve their capabilities in UI automation tasks.
Table 8. The comparison of task completion rate between mobile agents. Tasks are categorized to different difficulty levels according to the number of steps required to finished it in the ground-truth dataset.
@@ 5. Limitations of LlamaTouch
Supporting WebView-based apps. One major limitation of LlamaTouch is that its evaluation process requires Android’s view hierarchy to empower approximate UI layout matching and accurate UI component matching ($\S$§3.2). While some Android applications built with WebView (Google, 2024a) are unable to access their view hierarchy, making LlamaTouch cannot evaluate agent performance on these applications. However, only a minor number of applications are built with WebView, inspiring LlamaTouch’s solution in utilizing view hierarchy. When evaluating WebView-based apps are required, more advantaged techniques, such as screen similarity detection (Wu et al., 2023b, a), OCR/model-powered screen element recognition (Bunian et al., 2021; Xie et al., 2020), should be incorporated to retrofit LlamaTouch. We will consider this in future work.
Biases in identifying essential states. LlamaTouch requires human or other autonomous agents, e.g., GPT-4V, to identify and annotate essential states on pre-defined task execution traces. This process, however, could involve biases and leads to inaccuracy due to potentially limited knowledge in application execution and unseen circumstances. For example, given a task “Delete YouTube in the Google Play Store”, the pre-defined traces might explicitly navigate to the app page of “YouTube” and click the “Uninstall” button, which might be annotated as one essential state during task execution. However, they may ignore that this application might be not installed at all. In this case, there will be no explicit action of “clicking the uninstall button” while this task is still completed. Such biases might result in low accuracy in checking task completion rates. However, we think this limitation could be eliminated and refined by involving expert reviewing. Furthermore, for every single task description, involving diverse essential states on different potential task execution paths could better enhance the robustness of LlamaTouch’s evaluation design.
@@ 6. Conclusion
In this work, we proposed LlamaTouch, the first testbed for evaluating mobile agents with faithfulness and scalability in UI automation tasks. LlamaTouch enables mobile agents to be tested on real mobile devices. At the evaluation stage, it only matches the task execution traces with annotated essential states. LlamaTouch well tolerates different task execution paths and dynamic execution environments, significantly reducing false negative results that occurred in previous evaluation approaches. It achieves high evaluation accuracy like human validation, while preserving the scalability of evaluating on static datasets. By putting task execution on real mobile devices, we also reveal the limited capabilities of current mobile agents in UI automation tasks.
@@ Appendix A Appendix
A.1. APIs Provided by LlamaTouch

### v2

@@ 5.1. Experiment Setup
Mobile environments. LlamaTouch primarily utilizes an x86-64 Android emulator (Google, 2023) as the mobile environment for task execution. The Android emulator is configured with Android 12 (API level 31) with Google Play Services deploying on a Linux server running Ubuntu 18.04 OS. Tasks involving applications that are not compatible with the Android emulator (e.g., Snapchat) are tested on a real Google Pixel 5 device with Android 14 OS.
Task setup. In this study, tasks are set up in two stages. First, we construct an Android system image for the Android emulator, which provides a concrete environment where partial task states are prepared. This ensures a controlled and replicable starting point for the tasks. Second, we develop setup scripts using the Android UIAutomator2 library (uia, 2024) to manage complex setup processes with specific constraints (e.g., require manual login or run on real-world smartphones). These scripts automate the initialization of the environment, ensuring tasks are ready for execution with minimal manual intervention.
@@
  • Auto-UI (Zhan and Zhang, 2023) employs a multimodal transformer model based on BLIP-2 (Li et al., 2023) and T5 (Raffel et al., 2020) variants as the brain for decision-making. It takes pixel-level screenshots and task descriptions as input.
  • AutoDroid (Wen et al., 2024) uses LLMs for device control. It first takes textual screen VH and simplifies it to an explicit, readable HTML representation. The simplified HTML representation is then processed by LLMs to generate the corresponding action. We selected GPT-4 (gpt-4-0125-preview) (Achiam et al., 2023) as the backend of AutoDroid.
  • AppAgent (Yang et al., 2023) utilizes GPT-4V (OpenAI, 2023) to comprehend the screenshots of mobile devices and then dispatch controlling commands to LlamaTouch. In our evaluation, we used the AppAgent version without document pre-exploration for simplicity.
  • CoCo-Agent (Ma et al., 2024) uses LLaVA (Liu et al., 2023) as the brain. It is trained on a subset of the original AITW dataset.
@@
  • Step-wise action match is widely-utilized in evaluating mobile agents on well-established datasets (Rawles et al., 2023; Sun et al., 2022; Wen et al., 2024; Zhan and Zhang, 2023). It compares the agent-generated action sequences in real-world environments with ground-truth action sequences in the dataset. When two action sequences are identical, the task is treated as completed. We utilize the action match algorithm in AITW (Rawles et al., 2023): two actions are matched only if they have identical action types and parameters.
@@
5.2. Metrics
We primarily compare (1) the end-to-end task completion rate (TCR) of different mobile agents and (2) the accuracy of different evaluation approaches, using human validation results as the ground truth.
End-to-end TCR. Executing tasks in realistic mobile environments results in two outcomes: task completion and task non-completion. In step-wise action match, a task is considered completed only when the two action sequences are identical. LCS-based action match assesses whether a ground-truth action sequence is a subsequence of the agent-generated action sequence; if so, the task is evaluated as completed. In LlamaTouch, a task is considered completed when the agent execution trace passes through all essential states in the ground-truth dataset, using the proposed algorithm in §$\S$3.3. The end-to-end TCR of a certain evaluation method is calculated as the number of tasks evaluated completed divided by the total number of tasks in the dataset.
Evaluation accuracy. In this study, we use human validation results on agent execution traces as the ground truth for task completion. Assume $N$ denotes the ground-truth dataset encompassing all tasks. For task $n$ in $N$, we use $H_{n}$ to represent the human validation result based on the agent execution trace, where $H_{n}$ can be either “True” or “False”. When evaluating the same agent execution trace using a specific evaluation method, the outcome is represented by $E_{n}$, which can also be “True” or “False”. Therefore, the accuracy formula is defined as:
(1) $\text{accuracy}=\frac{\sum_{n\in N}\left(\delta_{E_{n},H_{n}}\right)}{|N|}$
The formulation indicates that a task is accurately evaluated only when the evaluation result $E_{n}$ matches the human validation result $H_{n}$. Consequently, the accuracy of an evaluation approach is defined as the proportion of correctly evaluated tasks to the total number of tasks in the dataset $|N|$.
@@ 5.3. Task Completion Rate and Accuracy
To demonstrate LlamaTouch’s effectiveness in addressing the issue of false negative results pervasive in previous evaluation methods, we focused on tasks considered successfully completed by human evaluation, excluding all incomplete tasks. Table 7 displays the number of tasks completed by agents and the evaluation accuracy for these tasks. Overall, agents successfully completed 30 tasks on average. Among these tasks, both step-wise action match and LCS-based action match achieve no more than 0.1% evaluation accuracy. This indicates they are unable to faithfully evaluate tasks executed in real-world environments using static datasets, primarily because a single static action sequence is difficult to match with multiple possible paths to task completion in real-world environments. Notably, LlamaTouch exhibits an average accuracy of 79% in validating these task execution traces, significantly reducing the percentage of false negative cases observed with other evaluation methods. The findings uncover LlamaTouch’s proficiency in evaluating UI automation tasks under real-world settings.
@@
5.4. Ablation Study
LlamaTouch achieves high accuracy when evaluating agent execution traces in real-world environments, attributed to the annotation and implementation of various essential state primitives. In this section, we evaluate the effectiveness of two types of match designs in LlamaTouch: fuzzy match and exact match. We aggregate the execution traces of all mobile agents evaluated in §$\S$5.3. We evaluate the system by first disabling exact match and fuzzy match separately, and then enabling different primitives one at a time. We present the results for (1) all tasks, (2) tasks in AITW (Rawles et al., 2023), and (3) new tasks generated in LlamaTouch. The results for task completion rate and accuracy are shown in Table 8.
Exact match ensures key information on two screens is identical. As shown in the results, exact match significantly contributes to improving evaluation accuracy. For example, without all exact match primitives, LlamaTouch’s evaluation accuracy on all tasks drops from 95% to 19%. Among all exact match primitives, activity match greatly improves evaluation accuracy. This is due to its simple yet efficient ability to locate functional application screens and the large proportion of annotated activity primitives, comprising 35% of all primitives in our dataset. UI component match is also necessary for faithful evaluation in LlamaTouch, as it is typically with screen location primitives such as activity to detect critical information within the matched screen. Exact match for action and system state only improves accuracy slightly because only a few tasks are annotated with these primitives.
Fuzzy match does not show a notable improvement in evaluation accuracy, despite the presence of 30 screen-level fuzzy match primitives and 52 textbox fuzzy match primitives in the dataset. Few tasks with these primitives can be completed by the four agents we evaluated, leading to only a slight improvement. Although primitives for fuzzy match are not well explored, they play an important role in dealing with screen content or UI layout dynamics in the real world. With more performant mobile agents in the future, they will be further explored and evaluated.
@@ 5.4. Ablation Study
Evaluation design All tasks AITW Generated
@@ 5.5. Absolute Capabilities of Mobile Agents
In this section, we present the absolute capabilities of different mobile agents in mobile UI automation tasks. First, we categorize tasks according to their sources: AITW and our self-generated dataset. This aims to determine whether an agent that has previously learned on a well-established dataset can adapt to new applications/tasks. Second, we categorize tasks according to their complexities, measured by the number of steps required to complete a task in the ground-truth dataset. Tasks are further divided into three difficulty levels: easy (steps $\leq$ 4), medium (4 $<$ steps $\leq$ 8), and high (step $>$ 8).
Performance across datasets. Among the four evaluated agents, AutoUI and CoCo-Agent were previously trained on AITW. AutoDroid and AppAgent directly invoke GPT-4 and GPT-4V, respectively. We separately show the end-to-end TCR of all agents on tasks in AITW and LlamaTouch’s generated tasks to demonstrate their performance generalization to new scenarios (i.e., new tasks and new applications). Results in Table 9 indicate that AutoDroid and AppAgent also achieve higher TCR in AITW tasks compared to LlamaTouch’s generated tasks. This is mainly attributed to the task complexity gap between the two datasets. AutoUI and CoCo-Agent achieve a 12.75% TCR on tasks in AITW. However, for generated tasks in LlamaTouch, both perform poorly, with only 2.29% and 2.31% TCR, respectively. The TCR gap between datasets reveals their lack of capability to adapt to previously unseen tasks. Considering the vast number of real-world applications, mobile agents with strong generalization capabilities to unseen scenarios are more competitive.
Table 9. The comparison of task completion rate categorized by task sources (i.e., from AITW and LlamaTouch generated tasks).
@@
Performance under different task complexity. We categorize tasks into three difficulty levels based on the number of steps required to complete them in the datasets. The results are shown in Table 10. Generally, mobile agents can better complete simple tasks that require fewer steps. There is a significant drop in TCR when task complexity increases. AutoDroid outperforms the other three agents for tasks at all difficulty levels. AppAgent achieves a TCR similar to AutoDroid across all tasks. We believe that the knowledgeable GPT-4/4V can better interpret task descriptions in natural language and pixel-level or textual UI representations. Therefore, agents based on GPT-4/4V achieve high TCR. However, there is still considerable room for mobile agents to improve their capabilities in mobile UI task automation.
Table 10. The comparison of task completion rate between mobile agents. Tasks are categorized into different difficulty levels according to the number of steps required to finish it in the ground-truth dataset.
@@ 6. Limitations of LlamaTouch
Supporting WebView-based apps. One limitation of LlamaTouch is that its evaluation process requires Android’s VH to empower approximate UI layout matching and accurate UI component matching (§$\S$3.2). Some Android applications built with WebView (Google, 2024a) are unable to access their VH, preventing LlamaTouch from evaluating agent performance on these applications. However, only a minor number of applications are built with WebView, inspiring LlamaTouch’s solution in utilizing VH. When evaluating WebView-based apps are required, more advantaged techniques, such as screen similarity detection (Wu et al., 2023b, a), OCR/model-powered screen element recognition (Bunian et al., 2021; Xie et al., 2020), should be incorporated to retrofit LlamaTouch. We will consider this in future work.
Biases in identifying essential states. LlamaTouch requires human or other autonomous agents, e.g., GPT-4V, to identify and annotate essential states on predefined UI interaction traces. This process, however, could involve biases and leads to inaccuracy due to potentially limited knowledge in application execution and unseen circumstances. For example, given a task “Delete YouTube in the Google Play Store”, the predefined traces might explicitly navigate to the app page of “YouTube” and click the “Uninstall” button, which might be annotated as one essential state during task execution. However, they may ignore that this application might be not installed at all. In this case, there will be no explicit action of “clicking the uninstall button” while this task is still completed. Such biases might result in low accuracy in checking task completion rates. However, we think this limitation could be eliminated and refined by involving expert reviewing. Furthermore, for every single task description, involving diverse essential states on different potential task execution paths could better enhance the robustness of LlamaTouch’s evaluation design.
@@ 7. Conclusion
In this work, we proposed LlamaTouch, the first testbed for evaluating mobile agents with both faithfulness and scalability in mobile UI task automation. LlamaTouch enables mobile agents to be tested on realistic mobile environments. At the evaluation stage, it matches the task execution traces with annotated essential states. LlamaTouch tolerates different task execution paths and dynamic execution environments, significantly reducing false negative results that occurred in previous evaluation approaches. It achieves high evaluation accuracy, which is comparable to human validation, while preserving the scalability of evaluating on static datasets. By conducting task execution to real-world mobile devices, we also reveal the limited capabilities of current mobile agents in mobile UI automation tasks.
Acknowledgements.
This work was supported by National Key R&D Program of China (No.2021ZD0113001), NSFC (No.62102045 and No.62272261), and China Institute of IoT (Wuxi).
@@ Appendix A Appendix
A.1. APIs Provided by AgentEnv